# Antichiral edge states and Bogoliubov Fermi surfaces in a two-dimensional proximity-induced superconductor

### v2

@@ I.1 First order in the SOC
¯$\overline{\cal H}({\bf k})$ $=$ $\frac{1}{2}\left[\left(\xi_{{\bf k}}-V\right)c^{\dagger}_{{\bf k},+}c_{{\bf k},+}+\left(\xi_{{\bf k}}+V\right)c^{\dagger}_{-{\bf k},-}c_{-{\bf k},-}\right]-\frac{\Delta_{0}}{2}\left(c^{\dagger}_{{\bf k},+}c^{\dagger}_{-{\bf k},-}+{\rm h.c.}\right).$ (S1)
@@
I.2 Linear order in the pairing, linear order in the SOC
Another route to derive the effective Hamiltonian Eq. (I.1) is to start by diagonalizing the normal part of the Hamiltonian including the kinetic, the SOC and the Zeeman fields.
@@ I.2 Linear order in the pairing, linear order in the SOC
¯$\overline{\cal H}_{\rm eff}({\bf k})\simeq\left(\xi_{\bf k}-V-\alpha_{R}k_{y}\right)c_{{\bf k},-}^{\dagger}c_{{\bf k},-}-\left(i\frac{\Delta_{0}\alpha_{R}k_{x}}{2V}c^{\dagger}_{{\bf k},-}c^{\dagger}_{-{\bf k},-}+{\rm h.c.}\right),$ (S14)
@@
Recalling that the topological phase takes place for $\mu^{2}\leq V^{2}-\Delta_{0}^{2}$, we see that Eq. (S15) agrees with Eq. (I.1) up to corrections $\propto\Delta_{0}^{2}$ within the range of parameters defining this phase.
II Calculation of the edge states in the effective continuum model
@@
III Details on the calculation of the Josephson current
@@
The system consists of two superconductors, $L$ and $R$, which are semi-infinite along the $x$- direction and modeled by a tight-binding Hamiltonian. The interface between the two superconductors is represented by a single site (see sketch of Fig.4 in the main text). Periodic boundary conditions are assumed along the $y$ direction, giving rise to $N_{y}$ transverse channels labeled by $k_{y}$.
@@
+$^{\dagger}_{\alpha,l_{x},k_{y}}\left[\tau^{z}(-t-is_{\alpha}\lambda\sigma^{y})\right]{\bf c}_{\alpha,l_{x}+1,k_{y}}+\text{h.c.}\right],$.c. ] ,
@@
[…]
@@
being ${\bf g}^{r}_{\alpha}$ ($\alpha=L,R$) the surface Green’s function of the superconductor at the site adjacent to the interface. This is computed using the same recursive algorithm as in Ref. [58, 24]. The advanced Green’s functions are obtained via ${\bf G}_{ij}^{a}(k_{y},\varepsilon)=\left[{\bf G}^{r}_{ji}(k_{y},\varepsilon)\right]^{\dagger}$.
@@
[…]

### v3

@@ II.1 First order in the SOC
$\overline{\cal H}({\bf k})$ $=$ $\frac{1}{2}\left[\left(\xi_{{\bf k}}-V\right)c^{\dagger}_{{\bf k},+}c_{{\bf k},+}+\left(\xi_{{\bf k}}+V\right)c^{\dagger}_{-{\bf k},-}c_{-{\bf k},-}\right]-\frac{\Delta_{0}}{2}\left(c^{\dagger}_{{\bf k},+}c^{\dagger}_{-{\bf k},-}+{\rm h.c.}\right).$ (S1)
@@
II.2 Linear order in the pairing, linear order in the SOC
Another route to derive the effective Hamiltonian Eq. (II.1) is to start by diagonalizing the normal part of the Hamiltonian including the kinetic, the SOC and the Zeeman fields.
@@ II.2 Linear order in the pairing, linear order in the SOC
$\overline{\cal H}_{\rm eff}({\bf k})\simeq\left(\xi_{\bf k}-V-\alpha_{R}k_{y}\right)c_{{\bf k},-}^{\dagger}c_{{\bf k},-}-\left(i\frac{\Delta_{0}\alpha_{R}k_{x}}{2V}c^{\dagger}_{{\bf k},-}c^{\dagger}_{-{\bf k},-}+{\rm h.c.}\right),$ (S14)
@@
Recalling that the topological phase takes place for $\mu^{2}\leq V^{2}-\Delta_{0}^{2}$, we see that Eq. (S15) agrees with Eq. (II.1) up to corrections $\propto\Delta_{0}^{2}$ within the range of parameters defining this phase.
III Calculation of the edge states in the effective continuum model
@@
IV Details on the calculation of the Josephson current
@@
The system consists of two superconductors, $L$ and $R$, which are semi-infinite along the $x$- direction and modeled by a tight-binding Hamiltonian. The interface between the two superconductors is represented by a row of sites (see sketch of Fig.4 in the main text). Periodic boundary conditions are assumed along the $y$ direction, giving rise to $N_{y}$ transverse channels labeled by $k_{y}$.
@@
$\left.+{\bf c}^{\dagger}_{\alpha,l_{x},k_{y}}\left[\tau^{z}(-t-is_{\alpha}\lambda\sigma^{y})\right]{\bf c}_{\alpha,l_{x}+1,k_{y}}+\text{h.c.}\right],$
@@
$J(\phi)=\frac{e}{\hbar}\sum_{k_{y}}\mbox{Re}\left\{\mbox{Tr}\left[\tau^{z}\sigma^{0}\;\hat{\tau}(\phi){\bf G}^{<}_{d,L}(k_{y};t,t)\right]\right\}=\frac{e}{h}\sum_{k_{y}}\mbox{Re}\left\{\int d\varepsilon\mbox{Tr}\left[\tau^{z}\sigma^{0}\;\hat{\tau}(\phi){\bf G}^{<}_{d,L}(k_{y},\varepsilon)\right]\right\},$ (S25)
@@
being ${\bf g}^{r}_{\alpha}$ ($\alpha=L,R$) the surface Green’s function of the superconductor at the site adjacent to the interface. This is computed using the same recursive algorithm as in Ref. [Gruneiro2023Jul, Mateos2024Aug]. The advanced Green’s functions are obtained via ${\bf G}_{ij}^{a}(k_{y},\varepsilon)=\left[{\bf G}^{r}_{ji}(k_{y},\varepsilon)\right]^{\dagger}$.
@@
$J(\phi)=\frac{e}{h}\sum_{k_{y}}\left\{\int d\varepsilon f(\varepsilon)F_{k_{y}}(\varepsilon)\right\},$ (S34)